Driving down a mountain, Tom finds that he has descended 1800 ft in elevation by the time he is 3.25 miles horizontally away from the top of the mountain. Find the slope of his descent to the nearest hundredth.
Answer by stanbon(75887)

Driving down a mountain, Tom finds that he has descended 1800 ft in elevation by the time he is 3.25 miles horizontally away from the top of the mountain. Find the slope of his descent to the nearest hundredth.
--------------
1 mi = 5280 ft.
1800 ft (1/5280) = 0.34 miles
-------
slope = (change in elevation)/(horizontal change) = 0.34/3.25= 0.105
==========
Cheers,
Stan H.